OpenStudy (pphalke):

graph the equation and find the y intercept x+y=-2

OpenStudy (pphalke):

0+y=2?

OpenStudy (anonymous):

x + y = -2 -x -x y = -x - 2

OpenStudy (pphalke):

what do I do now @larryboxaplenty

OpenStudy (dayakar):

first u have to draw a line to draw a line minimum two points are required

OpenStudy (anonymous):

You don't need to draw a line, you can solve via shortcut

OpenStudy (pphalke):

how do I solve for the y intercept?

OpenStudy (anonymous):

y = -x-2, let x = 0 to find y-intercept, let y=0 to find x-intercept.

OpenStudy (pphalke):

so would it be y=-0-2??

OpenStudy (pphalke):

is the answer y=-0-2=-2 check plz @larryboxaplenty

OpenStudy (unklerhaukus):

yes, the y-intercept is when x=0